OpenStudy (anonymous):

Given the following function f(x)=√3 sinx+cosx on the interval [0,2π]. A. Determine the open interval(s) where f(x) is increasing. B. Determine the open interval(s) where f(x) is decreasing. C. Determine the relative maximum(s) or minimum(s).

OpenStudy (anonymous):

\[\text{Re-write }\sqrt{3} \sin x + \cos x\ \text{ in the form } R\sin (x + \alpha)\] ... and go from there.

OpenStudy (anonymous):

???...showing math processing error

OpenStudy (anonymous):

\[\sqrt{3}\sin x + \cos x = 2\sin\left(x+\frac{\pi}{6}\right)\]

OpenStudy (anonymous):

That probably makes it easier to work out what you need from? I would explain how to get that normally, but very tired ¬_¬.